OpenStudy (anonymous):

At a certain electronics company, the daily output Q is related to the # of people on the assembly line by Q = 400 + 10 sqrt(A + 25) (a) determine output if there are 44 people on the assembly line. (b) determine how many people are needed on the assembly line if the required daily output is 520.

OpenStudy (anonymous):

\[Q = 400 + 10 \sqrt{(A + 25)}\] a) A=44, so: \[Q = 400 + 10 \sqrt{(44 + 25)} = 400 + 10\sqrt{69}\approx483.07\] b) Q = 520, so: \[520 = 400 + 10 \sqrt{(A + 25)}\] \[120= 10 \sqrt{(A + 25)}\] \[12= \sqrt{(A + 25)}\] \[A + 25 = 144\] \[A = 144 -25=119\]